Given that M, A, T, and H are positive integers where 9T = 2H, 5M = 3A, and
10H = 9A. Arrange M, A, T, and H in order from least to greatest.

Answer by Fombitz(32388)

9T=2H
5M=3A
10H=9A
H=(9/10)A
.
.
9T=2H
T=(2/9)H=(2/9)(9/10)A=(1/5)A
.
.
5M=3A
M=(3/5)A
.
.
T, M, H, A